# A Graph Theoretic Approach for Exploring the Relationship between EV Adoption and Charging Infrastructure Growth

### v1

@@ Abstract
The increasing global demand for conventional energy has led to significant challenges, particularly due to rising CO2 emissions and the depletion of natural resources. In the U.S., light-duty vehicles contribute significantly to transportation sector emissions, prompting a global shift toward electrified vehicles (EVs). Among the challenges that thwart the widespread adoption of EVs is the insufficient charging infrastructure (CI). This study focuses on exploring the complex relationship between EV adoption and CI growth. Employing a graph theoretic approach, we propose a graph model to analyze correlations between EV adoption and CI growth across 137 counties in six states. We examine how different time granularities impact these correlations in two distinct scenarios: Early Adoption and Late Adoption. Further, we conduct causality tests to assess the directional relationship between EV adoption and CI growth in both scenarios. Our main findings reveal that analysis using lower levels of time granularity result in more homogeneous clusters, with notable differences between clusters in EV adoption and those in CI growth. Additionally, we identify causal relationships between EV adoption and CI growth in 137 counties, and show that causality is observed more frequently in Early Adoption scenarios than in Late Adoption ones. However, the causal effects in Early Adoption are slower than those in Late Adoption.
@@ 1 Introduction
Achieving zero carbon emissions in light-duty transportation is crucial, and electric vehicles (EVs) offer a promising solution, as they produce zero tailpipe emissions [4]. However, widespread EV adoption faces obstacles, such as insufficient charging infrastructure, high upfront costs, and concerns over limited range [5]. Government interventions, including subsidies for charging infrastructure (CI) and tax credits for EV purchases, aim to address these barriers. Despite these efforts, the relationship between EV adoption and CI development remains complex: does the availability of charging stations drive EV adoption, or does widespread EV use lead to more charging infrastructure? Furthermore, the impacts of one on the other may emerge over different time granularities.
@@
This study explores the relationship between EV adoption (represented by the EV growth rate) and CI growth using a graph-theoretic approach to propose a graph model for analyzing correlations in EV adoption across 137 counties in six U.S. states. In the proposed graph model, nodes represent counties, and edges represent correlations between them. Our analysis constructs a graph model to capture correlations at various time granularities: monthly, bi-monthly, quarterly, and bi-annual. After constructing a graph for each time granularity, we apply the Louvain method, a community detection technique, to cluster counties into different groups. Similarly, we perform the same analysis on correlations in CI growth across these counties. We then examine the relationship between EV adoption and CI growth considering different time lags to explore Early Adoption scenarios (where EV adoption leads CI growth) and Late Adoption scenarios (where CI growth leads EV adoption). Finally, we analyze causal relationships between EV adoption and CI growth by considering various combinations of time granularities and lags to better understand these complex interactions.
@@
Firstly, our analysis reveals that, within EV adoption networks, counties from the same state tend to cluster together, exhibiting greater homogeneity at finer time granularities. However, this pattern is not observed in CI growth networks. Secondly, the correlations between EV and CI networks in both adoption scenarios, Early Adoption and Late Adoption, are consistently weakly positive, with Late Adoption correlations being slightly higher than those in Early Adoption. Thirdly, 115 out of 137 counties exhibit causal relationships between EV adoption and CI growth, analyzed using various combinations of time granularities, time lags, and the two adoption scenarios. Notably, causality is observed more frequently in Early Adoption scenarios than in Late Adoption ones. However, the causal effects in Early Adoption are slower than those in Late Adoption, which can be attributed to the significantly higher costs of establishing new CIs compared to adopting new EVs.
@@ 2.1 Impact of Charging Infrastructure and Other Factors on EV Adoption
Ledna et al. investigated the effect of various support policies for infrastructure investment and vehicle purchases on EV adoption in California [6]. They analyzed three policy scenarios: no support, support for either infrastructure or vehicle purchases, and combined support for both. These scenarios were assessed under conservative and optimistic technological advancement assumptions. Their findings indicated that the combined policy best supported vehicle sales and reduced CO2 emissions, while support focused on infrastructure alone proved more effective in conservative scenarios.
In another study, Chen et al. proposed a game-theoretic model to understand the interactions between the government, technology firms, and consumers in the EV market [5]. Their objective was to identify optimal government incentives to maximize market penetration while enhancing consumer surplus and social welfare. Their research revealed that, under a high-budget incentive framework, usage subsidies best maximized welfare, while investment subsidies were optimal for market penetration.
White et al. examined the relationship between the intent to adopt EVs and charging infrastructure through mediation analyses in three U.S. metropolitan areas [7]. They evaluated the indirect effects of Range Anxiety, Perceived Mobility Restriction (PMR), and Subjective Norms on adoption intent. Their results demonstrated that Subjective Norms were the strongest mediator, while Range Anxiety was marginally significant, and PMR showed no significant effect.
Khan et al. hypothesized that socio-demographic features and highway routes influence the presence of charging stations at the zip-code level [8]. They compared zip codes with and without stations using a t-test, finding that the presence of highways, income, and education levels positively correlated with charging station distribution.
At a country level, in Australia, Broadbent et al. proposed a framework to estimate the number of EVs in parking lots in Melbourne and transform them into EV-enabled spaces [9]. In Germany, Illmann et al. analyzed monthly data from 2011 to 2017 and found a long-term relationship between registered EVs and charging stations [10]. In France, Haidar et al. discovered that different incentives impacted hybrid and fully electric vehicles differently [11].
@@ 2.2 Application of Graph Model to Charging Infrastructure
The graph model is a powerful tool for analyzing relationships between various entities. A correlation network, where nodes represent entities or variables, and edges indicate correlation strength, is widely used in diverse fields, including bioinformatics, engineering, and social sciences.
The application of the graph model to the charging infrastructure has focused primarily on optimizing the distribution of charging stations [12]. For example, Gagarin et al. applied the dominating set concept to create a network of charging stations reachable within a specific threshold [13]. Similarly, Arkin et al. utilized Unit Disk Graphs to determine the minimum independent set required to cover an area with charging stations [14]. Moreover, studies in [15, 16, 17] applied the concept of the hitting set to identify the optimal number and locations for charging stations. Finally, Altundogan et al. combined genetic algorithms and shortest path optimization to determine the best locations for a pre-determined number of charging stations [18].
@@ 3.2 Proposed Graph Theoretic approach: correlation network
Our methodology applies modularity optimization to determine the optimal correlation threshold for each correlation network. The threshold that yields the highest modularity is selected as the most suitable for that network. This approach ensures that the network structure, defined by the retained correlations, reveals the most meaningful and interpretable relationships. Subsequently, the Louvain method is employed to detect communities (clusters) within the network.
@@ 3.4 Different Time Granularities
The relationship between EV adoption and CI growth is complex; therefore, our analysis considers using various time granularities to capture hidden relationships that may emerge at different time scales. To this end, we divide the dataset into multiple time granularities: monthly, bi-monthly, quarterly, and bi-annual periods. For each granularity, we calculate the growth rate from one period to the next. For instance, in the monthly granularity, we compute the growth rate from January to February, February to March, and so on. For the quarterly granularity, we calculate growth from the first quarter to the second quarter, and so forth.
Through this analysis, we examine the correlation networks of EV adoption and the growth of various CI levels, observing these correlations across different time scales to uncover potential relationships between EV adoption and CI growth. Specifically, we investigate whether clusters observed at finer granularities persist at coarser levels or whether new correlations between counties emerge, thereby shedding light on the temporal dynamics of EV adoption and CI expansion.
@@ 4 Results
In this section, we present the outcomes of our analysis, including the correlations for both EV adoption and CI growth using Graph Theory. We examine the two cases: Early Adoption (Section 4.1) and Late Adoption (Section 4.2) using different time granularity. Additionally, we present the results of our analysis of causality relationships between EV adoption and CI growth in these two cases in (Section 4.3).
@@ 5 Conclusion and Future Directions
In this study, we examined the relationship between EV adoption and charging infrastructure (CI) growth using Graph Theory and causality analysis. Our findings highlight that, within the correlation networks of EV adoption, counties from the same state tend to cluster together, exhibiting greater homogeneity at finer time granularities. However, this pattern is not observed in CI growth networks. The homogeneity in clusters of counties for EV adoption suggests the importance of internal state-level characteristics that drive similar patterns of EV adoption within their counties.
Moreover, the correlations between EV adoption and CI growth are negligible in the two adoption scenarios: Early Adoption and Late Adoption. These correlations are very weakly positive, with slightly higher values in the latter scenario, where sufficient CI must be provided before EVs are widely adopted. This finding marginally favors the provision of sufficient CI as a strategy to promote EV adoption.
In our analysis of the causal effects between EV adoption and CI growth, we find a higher number of counties favoring Early Adoption compared to Late Adoption. This emphasizes the importance of policies and incentives aimed at boosting early EV adoption to spur subsequent investments in charging infrastructure. Additionally, our analysis of causal relationships shows that the causal effect of EV adoption on CI growth is slower than the reverse. Finally, 115 out of 137 counties demonstrate causal relationships under different combinations of time granularity, time lag, and adoption scenarios.
One potential direction for future work is to conduct an enrichment analysis of other factors, including demographic, economic, and political variables, for the counties that tend to cluster together. This would provide deeper insights into why these counties exhibit similar patterns in EV adoption. Another avenue for future research is to explore how different levels of CI relate to the two cases of Early and Late Adoption.

### v2

@@ Abstract
The increasing global demand for conventional energy has led to significant challenges, particularly due to rising CO2 emissions and the depletion of natural resources. In the U.S., light-duty vehicles contribute significantly to transportation sector emissions, prompting a global shift toward electrified vehicles (EVs). Among the challenges that thwart the widespread adoption of EVs is the insufficient charging infrastructure (CI). This study focuses on exploring the complex relationship between EV adoption and CI growth. Employing a graph theoretic approach, we propose a graph model to analyze correlations between EV adoption and CI growth across 137 counties in six states. We examine how different time granularities impact these correlations in two distinct scenarios: Early Adoption and Late Adoption. Further, we conduct causality tests to assess the directional relationship between EV adoption and CI growth in both scenarios. Our main findings reveal that analysis using lower levels of time granularity result in more homogeneous clusters, with notable differences between clusters in EV adoption and those in CI growth. Additionally, we identify causal relationships between EV adoption and CI growth in 137 counties, and show that causality is observed more frequently in Early Adoption scenarios than in Late Adoption ones. However, the causal effects in Early Adoption are slower than those in Late Adoption
@@ 1 Introduction
Achieving zero carbon emissions in light-duty transportation is a crucial step toward this goal, and electric vehicles (EVs) offer a promising solution, as they produce zero tailpipe emissions [4]. However, widespread EV adoption faces several challenges, including insufficient charging infrastructure, high upfront costs, and concerns about limited driving range [5]. To mitigate these barriers, government interventions have been implemented, such as subsidies for charging infrastructure (CI) and tax credits for EV purchases. Therefore, a sufficient charging infrastructure is significant to sustain an effective transition to EV [6].
Moreover, several studies have investigated the influence of charging infrastructure alongside other factors, such as socio-demographic, economic, environmental, and political variables, on EV adoption in the United States [7, 8, 9, 10]. While charging infrastructure consistently emerges as a significant driver of EV adoption, the sole relationship between EV adoption and charging infrastructure growth remains underexplored. Existing research typically embeds charging infrastructure within broader multivariable models, making it difficult to isolate its unique effect.
Thus, a key question remains open: Does widespread EV adoption stimulate the development of more charging infrastructure (Early Adoption), or does the availability of charging infrastructure drive EV adoption (Late Adoption)? Additionally, such relationships may vary depending on the temporal granularity considered. Hence, this paper contributes to the field by investigating the sole relationship between EV adoption rate and CI growth in two different scenarios of Early Adoption and Late Adoption (as defined in Section 3.3) through two complementary analyzes: (1) correlation analysis using a graph-theoretic approach and (2) causality analysis.
@@
In the correlation analysis, we propose a graph model for analyzing EV adoption patterns across 137 counties in six U.S. states, where nodes represent counties and edges represent correlations between their EV adoption rates. In each adoption scenario, we construct graph models at multiple time granularities, monthly, bi-monthly, quarterly, and bi-annual to capture temporal dynamics. Using the Louvain method, a community detection technique, we cluster counties into groups based on their correlation structures. A similar graph-theoretic analysis is performed on CI growth patterns. Subsequently, we compute the correlation between the EV adoption and CI growth at different combinations of adoption scenarios, and time granularities. In the causality analysis, we examine the directional causality relationship between EV adoption and CI growth for each county incorporating different time lags, and using the same combinations of adoption scenarios, and time granularities.
@@
Firstly, the study analysis reveals that, within EV adoption networks, counties from the same state tend to cluster together, exhibiting greater homogeneity at finer time granularities. However, this pattern is not observed in CI growth networks. Secondly, the correlations between EV and CI networks in both adoption scenarios, Early Adoption and Late Adoption, are consistently weakly positive, with Late Adoption correlations being slightly higher than those in Early Adoption. Thirdly, 115 out of 137 counties exhibit causal relationships between EV adoption and CI growth, analyzed using various combinations of time granularities, time lags, and the two adoption scenarios. Notably, causality is observed more frequently in Early Adoption scenarios than in Late Adoption ones. However, the causal effects in Early Adoption are slower than those in Late Adoption, which can be attributed to the significantly higher costs of establishing new CIs compared to adopting new EVs.
@@ 2.1 Impact of Charging Infrastructure and Other Factors on EV Adoption
There is an extensive body of literature examining electric vehicle (EV) adoption globally. Since our study focuses on the United States, we narrow our review to studies that are most relevant to our analysis—specifically, those conducted within the U.S. that consider charging infrastructure as a key factor influencing EV adoption. These studies provide the necessary background and support for our research approach. To underscore the broader significance of this topic, we also briefly highlight selected national-level studies from other countries at the end of this section.
@@
Ledna et al. investigated the impact of various support policies on EV adoption in California, focusing on investments in public charging infrastructure and vehicle purchase subsidies [11]. They simulated three policy scenarios: no support, support for either infrastructure or vehicle purchases, and combined support; under both conservative and optimistic assumptions regarding technological advancement. Their findings suggest that combined support policies are most effective in boosting EV sales and reducing CO2 emissions overall, while infrastructure-focused support yields better outcomes in conservative technology scenarios, emphasizing the importance of tailoring policy strategies to market and technological conditions.
In another study, Chen et al. developed a game-theoretic model to examine the strategic interactions among governments, technology firms, and consumers in promoting electric vehicle (EV) adoption under various subsidy policies [5]. Their goal was to identify optimal government incentive strategies, focusing on infrastructure investment subsidies versus usage subsidies, under constrained budgets. The study found that when the budget is limited, either type of subsidy can be effective. However, under a high-budget scenario, investment subsidies are more effective for maximizing market penetration, while usage subsidies better enhance consumer surplus and overall social welfare
White et al. examined the mechanisms underlying the relationship between public charging infrastructure and electric vehicle (EV) adoption intent across three major U.S. metropolitan areas: Los Angeles, Dallas/Fort Worth, and Atlanta [12]. Using multiple mediation analysis, they tested whether three psychological constructs—range anxiety, perceived mobility restriction (PMR), and subjective norms—mediated the effect of public charging station density on EV adoption intent. Their findings revealed that subjective norms were the strongest and most consistent mediator, significantly influencing adoption intent in all three regions. Range anxiety played a marginal role in two of the regions, while PMR showed no significant mediating effect. These results suggest that the visibility of charging infrastructure may function more as a social signal promoting adoption, rather than alleviating technical or mobility concerns.
Khan et al. examined the distribution of EV charging infrastructure across New York City to identify the socio-demographic and transportation factors associated with charging station access [13]. Using correlation and hypothesis testing at the zip-code level, they found that the presence of highways, higher median household income, and a greater percentage of White-identifying population were positively associated with the presence and number of charging stations. Their findings highlight significant disparities, with low-income and minority communities having notably less access to EV charging infrastructure, underscoring the need for equity-focused policies in infrastructure planning.
Burra et al. addressed the lack of sufficient EV-owning households in Maryland’s travel surveys by developing a synthetic population using a Bayesian network approach [7]. They estimated household-level EV ownership probabilities using a binomial logit model and evaluated how socio-demographic factors and access to Level 2 and DC fast charging infrastructure influence adoption. Their findings revealed that high-income and suburban households are more likely to own EVs, and that access to workplace charging and DC fast chargers significantly increases the likelihood of ownership. The study highlights the importance of differentiating between charger types and suggests that equitable infrastructure deployment should consider both location and income disparities.
Pamidimukkala et al. conducted a survey-based study of University of Texas at Arlington affiliates to examine how four categories of barriers—technological, environmental, financial, and infrastructure—affect consumers’ intentions to adopt electric vehicles (EVs) [8]. Using structural equation modeling on responses from 733 participants, they found that financial, infrastructure, and technological barriers all had significant negative effects on EV adoption intention, with financial barriers having the strongest influence. Infrastructure barriers—such as insufficient public charging stations and limited maintenance and repair services—also emerged as key obstacles. In contrast, environmental barriers were not statistically significant, suggesting that respondents largely accepted the environmental benefits of EVs.
Debnath et al. applied computational text analysis to approximately 36,000 Facebook posts—comprising a corpus of around 600,000 words—to explore public discourse surrounding electric vehicle (EV) adoption in the United States across six PESTLE dimensions: Political, Economic, Social, Technological, Legal, and Environmental [14]. Using a mixed-methods approach that combined social network analysis and machine learning-based topic modeling, they identified key themes within each category. The analysis revealed that discussions of charging technology and renewable energy dominated the Technological dimension, while concerns related to charging infrastructure—particularly public funding, regulation, and policy—were primarily clustered under the Political category.
Kamis et al. developed predictive models for three 2021 targets—EV registrations, EV-related jobs, and new public charging infrastructure (Level 2)—using county-level data across a wide range of predictors, including existing charging infrastructure, environmental conditions, education, and socio-demographic factors. They found that existing charging infrastructure was a significant predictor across all models, with Gradient Boosted Trees showing superior accuracy over traditional regression methods [9].
At the national level, Broadbent et al. developed a national-scale policy simulation model to explore how public incentives and infrastructure investment affect electric vehicle (EV) adoption trajectories and emissions reductions in Australia. Additionally, Babic et al. proposed a simulation-based operational model for transforming urban parking lots in Melbourne into EV-enabled charging hubs, focusing on profitability, charger allocation, and consumer behavior [15]. In Germany, Illmann et al. analyzed monthly ZIP-code-level data from 2012 to 2017 and found a significant long-term relationship between public charging infrastructure and private EV registrations, with fast chargers having the strongest influence [16]. In France, Haidar et al. used mixed-effects regression to analyze BEV and PHEV adoption across 94 departments and found that different incentives and charging infrastructure types affect each market differently, with fast and ultra-fast chargers driving BEV sales, and slow-to-normal chargers more relevant for PHEVs [17]. In China, studies examine the relationship between electric vehicle (EV) adoption and the availability of charging infrastructure. They consistently find that improvements in charging infrastructure have a positive impact on promoting EV adoption, though the nature of the effect depends on the policy design, regional conditions, and stakeholder interactions [18, 19].
@@ 2.2 Application of Graph Model to Charging Infrastructure
The graph model is a powerful tool for analyzing relationships between various entities. A correlation network, where nodes represent entities or variables, and edges indicate correlation strength, is widely used in diverse fields, including bioinformatics, engineering, and social sciences [20, 21, 22].
The application of graph-theoretic models to electric vehicle (EV) charging infrastructure has predominantly focused on optimizing station placement and network coverage. Gagarin et al. modeled the problem as a multiple domination problem on reachability graphs to ensure drivers can access multiple charging options within a limited cruising range [23]. Arkin et al. employed unit disk graphs and t-spanners to determine the minimum number of stations required to enable near-shortest EV routes with sufficient recharging support [24]. Moreover, studies in [25, 26, 27] applied the concept of the hitting set to identify the optimal number and locations for charging stations. Finally, Altundogan et al. integrated graph-based distance modeling with a genetic algorithm to optimize the spatial distribution of a fixed number of chargers across urban networks [28].
While these studies demonstrate the utility of graph theory for infrastructure planning, they do not explore the dynamic, temporal interplay between EV adoption trends and charging infrastructure growth—an area this study addresses by applying graph models to correlate and interpret adoption patterns directly.
@@ 3.2 Proposed Graph Theoretic approach: correlation network
To investigate the structural relationships among counties based on their electric vehicle (EV) adoption or charging infrastructure growth, we construct an unweighted, undirected graph where nodes represent counties and edges represent statistically significant correlations between them. Let $C_{i}$ and $C_{j}$ denote the time series associated with counties $i$ and $j$, respectively. The similarity between each pair of counties is measured using the Pearson correlation coefficient:
$\rho_{ij}=\frac{\mathrm{Cov}(C_{i},C_{j})}{\sigma_{C_{i}}\sigma_{C_{j}}},$
where $\mathrm{Cov}(C_{i},C_{j})$ is the covariance between the two time series, and $\sigma_{C_{i}}$ and $\sigma_{C_{j}}$ are their standard deviations. We construct a graph $G=(V,E)$, where each node $v\in V$ represents a county, and an edge $(i,j)\in E$ exists if $\rho_{ij}\geq\tau$, with $\tau$ being a correlation threshold selected through a modularity optimization process.
To identify community structure within the network, we apply the Louvain community detection algorithm, which partitions the graph into non-overlapping communities. The quality of the partition is assessed using the modularity metric, defined as:
$Q=\frac{1}{2m}\sum_{i,j}\left[A_{ij}-\frac{k_{i}k_{j}}{2m}\right]\delta(c_{i},c_{j}),$
where $A_{ij}$ is the adjacency matrix of the graph, $k_{i}$ and $k_{j}$ are the degrees of nodes $i$ and $j$, $m$ is the total number of edges, $c_{i}$ is the community assignment of node $i$, and $\delta(c_{i},c_{j})$ is the Kronecker delta function, equal to 1 if $c_{i}=c_{j}$ and 0 otherwise [31]. The correlation threshold $\tau$ is selected to maximize the modularity $Q$ and ensure meaningful community separation.
After constructing the initial graph and performing community detection, we apply a refinement procedure that iteratively removes the edge with the highest edge betweenness centrality, defined as the number of shortest paths passing through that edge. After each removal, the modularity is recalculated to evaluate improvements in community structure. This process helps expose more distinct and internally cohesive communities by removing structural bridges between densely connected groups of counties.
@@ 3.4 Different Time Granularities
The relationship between electric vehicle (EV) adoption and charging infrastructure (CI) growth is complex; therefore, our analysis explores multiple time granularities to capture underlying patterns that may emerge at different temporal resolutions. To this end, we divide the dataset into several time granularities: monthly, bi-monthly, quarterly, and bi-annual periods. For each granularity, we calculate the growth rate from one period to the next. For instance, at the monthly level, we compute the adoption or growth rate from January to February, February to March, and so on. At the quarterly level, we calculate growth from the first quarter to the second quarter, and so forth. The adoption or growth rate (AGR) at each step is mathematically expressed as:
$\mathrm{AGR}_{g_{j}}=\frac{g_{j}-\sum_{k=1}^{j-1}g_{k}}{\sum_{k=1}^{j-1}g_{k}}$
where $\mathrm{AGR}_{g_{j}}$ represents the growth rate at granularity $g_{j}$, $g_{j}$ is the number of registered EVs or newly established CI units during the current period, and $\sum_{k=1}^{j-1}g_{k}$ is the accumulated total from all previous periods.
Through this analysis, we examine the correlation networks of EV adoption and the growth of multiple CI levels (e.g., Level 1, Level 2, and DC fast charging) across different temporal granularities. Specifically, we investigate whether clusters identified at finer resolutions persist when observed at coarser levels, or whether new inter-county correlations emerge, thereby shedding light on the temporal dynamics of EV adoption and CI expansion across different types of infrastructure.
@@ 4 Results
In this section, we present the outcomes of our analysis, including the correlations for both EV adoption and CI growth using a graph theoretic approach. We examine the two cases: Early Adoption (Section 4.1) and Late Adoption (Section 4.2) using different time granularity. Additionally, we present the results of our analysis of causality relationships between EV adoption and CI growth in these two cases in (Section 4.3).
@@ 4.4 Outlier Counties in EV Adoption Clusters
5 Limitations and Future Work
This study focuses exclusively on the correlation and causal dynamics between electric vehicle (EV) adoption and the growth of charging infrastructure (CI), using a graph-theoretic model and time-series causality analysis. One inherent limitation is that the analysis does not explicitly isolate the influence of external factors such as policy interventions, technological advancement, and socio-demographic variables. Instead, we assume that such influences are implicitly embedded in the observed trends of EV adoption and CI growth across counties. For example, government incentives or socioeconomic conditions that encourage EV adoption are presumed to be reflected in the adoption data; likewise, incentives or budget allocations for infrastructure development are assumed to be manifested in the CI growth data. Although this assumption allows us to study the direct relationship between EVs and CIs, it restricts our ability to attribute causal mechanisms to specific policy or technological factors.
While our graph-theoretic framework reveals meaningful clustering patterns in both EV adoption and CI growth, it does not directly explain why certain counties tend to group together. Understanding the underlying drivers of these clusters—such as shared socio-economic profiles, policy environments, or infrastructure investments—remains an important next step. Future work should aim to uncover these commonalities by integrating additional contextual variables (e.g., population density, median income, political orientation, or state-level incentives) into the analysis. Whether through multivariate modeling or cluster enrichment techniques, such extensions would help determine whether the structural patterns observed in the networks reflect deeper systemic factors, thereby improving both interpretability and policy relevance. Furthermore, examining the role of different CI levels (e.g., Level 2 vs. DC fast chargers) in shaping adoption patterns under early and late adoption scenarios could offer a more granular understanding of the interplay between EV demand and infrastructure supply.
@@ 6 Conclusion and Future Directions
This study employs a graph-theoretic and time-series causality approach to investigate the relationship between EV adoption and charging infrastructure (CI) growth across 137 counties in six U.S. states. The results reveal consistent clustering patterns in EV adoption, particularly within states, whereas CI growth networks show weaker structural patterns. The correlation between EV and CI growth is generally weak across early and late adoption scenarios, with slightly stronger values observed when sufficient CI precedes EV adoptxion.
Our causality analysis indicates that while EV adoption does causally influence CI growth, the effect is slower compared to the reverse scenario, where CI growth more promptly triggers EV adoption. Importantly, over 80% of the counties analyzed exhibit significant causal relationships under at least one configuration of time lag, time granularity, and adoption scenario.
Despite the innovation in methodology, our analysis assumes that external drivers such as policy, economic incentives, and technology improvements are embedded within the temporal trends of the variables studied. Future work should incorporate multivariate analysis to isolate these effects and offer more actionable policy recommendations. Additionally, extending this framework to include county-level attributes may uncover structural explanations behind the observed clustering, enhancing our understanding of spatial and temporal patterns in EV adoption and CI expansion.